Josue had the following income expenses, gains and losses for 2019:

salaries received from big business, inc. $115,600

net business income from small business 49,500

net rental income from single family home 21,600

interest income 5,222

qualified dividend income 3,123

gain from sale of Alphabet common stock 12,200

How much is Josue's self-employment tax? show your work

How much is Josue's section 199A QBI deduction? show your work
Answer #1

Dear Student,

If you any doubts, feel free to ask

Part 1

Josue's self-employment tax

$3471

Self-employment income for self-employment tax = 49500*92.35% = 45713

Social Security wage limit is $132900 for 2019

The $17300 of $45713 will be subject to self-employment tax for OASDI and the total $45713 will be subject to self-employment tax for HI. No additional 0.9% medicare tax is charged as the total income of $161313 (115600+45713) subject to social security tax

Self-employment tax for OASDI (17300*12.4%)

2145

self-employment tax for HI (45713*2.9%

1326

Total self-employment tax

$3471

Part 2

Josue's section 199A QBI deduction

$9553

Josue's section 199A QBI deduction = (net income – (50%*self-employment tax))*20%

= (49500-(50%*3471))*20% = $9553